OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)?

OpenStudy (danjs):

radius is the distance from (0,0) to (-5,-7)

OpenStudy (danjs):

\[r = \sqrt{5^2 + 7^2} = \sqrt{74}\] center(h,k) = (-5,-7) Fill in (x-h)^2 + (y-k)^2 = r^2

OpenStudy (solomonzelman):



OpenStudy (danjs):

just putting in the numbers gives (x+5)^2 + (y+7)^2 = 74